Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (RounAssume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (RounAssume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (RounAssume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (RounAssume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Roun
Answer #1

Solution Given that u=48 6-16 P(40 < X<47) =P 40-48 < X-4 <47-48 =P [-0.5 <2 5-0.08] =P [zs -0.00 ] -p [25-0.5] = 0.4761 – 0.Date Page u=14.6 6=3.3 P(8 € 45 122 =P [ 8-14.6 8 < x - 14.6 3.3 -u < 12 -14.6 3.3 =P [-2 s 2 = -0.79] =P [ 25-0.79)-p [24-2]4= 38 6 = 14 PC 50 sxs 70) = P < 5-8 150-38 c 14. 4 1 3 < 70-38 14 =P [0.86 € 2 52.29] =P [ zs 2.2d] -p [250.86] = 0.9890 - 0classmate Date Page u=15.1 6= 3.8 PC 10 <x< 26) -P 10-15.1 LX-4 L 3. 8 6 526-15.1 3.8 -р . -5.1 3.8 10.9 3.8 =P [-1.34 <3 < 2